How large a sample is needed to estimate the population proportion for the following? ME = 0.03; α=0.05

a.) 962

b.) 1068

c.) 1297

d.) 1242
Answer #1

Solution :

Given that,

\hat p= 0.5

1 - \hat p = 1 - 0.5 = 0.5

margin of error = E = 0.03

At confidence level the z is ,

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.96 / 0.03)2 * 0.5 * 0.5

= 1068

Sample size =